Let f(x)=x^2 +3x+9 and g(x)=(2x+3)/(x-2) Find f(-1)-g(0)
Answer by nerdybill(7384)

Let f(x)=x^2 +3x+9 and g(x)=(2x+3)/(x-2) Find f(-1)-g(0)
.
first, find f(1):
f(x)=x^2 +3x+9
f(1)=(1)^2 +3(1)+9
f(1)=1+3+9
f(1)=13
.
then, find g(0):
g(x)=(2x+3)/(x-2)
g(x)=(2(0)+3)/(0-2)
g(x)=(3)/(-2)
g(x)= -3/2
.
finally:
f(-1)-g(0)
13-(-3/2)
13 + 3/2
26/2 + 3/2
(26+3)/2
29/2 (answer)